# TRAIL: Trust-Aware Client Scheduling for Semi-Decentralized Federated Learning

### v2

@@ 8             $x_{i}=1$;
From Theorem 1, the global training loss minimization initially outlined in the problem can be reinterpreted to focus primarily on minimizing the parameter $B$. This revised formulation, therefore, positions $B$ as the central target for reduction, aiming to directly influence and improve the overall system performance by addressing the underlying factors contributing to $B$ ’s value.
@@
To solve this nonlinear integer programming problem, we propose a greedy algorithm described in Algorithm 1.
@@
This section will evaluate our proposed mechanisms using four real-world datasets. First, we will introduce the basic setup of the experiments and the benchmark for comparison. Then, we will present the results of our experimental comparisons.
@@
  1. 1.
    Basic Setup:
    • We set up a system with five edge servers and fifty clients. Each client possesses 1,000 local training data samples. Additionally, we have set up the system so that 10%, 30%, and 50% of the clients experience a gradual decline in training quality and communication quality as the training progresses.
  2. 2.
    Datasets:
    • Real-world Datasets. Four standard real-world datasets, e.g. MNIST (Rana, Kabir, and Sobur 2023), EMNIST (Majeed et al. 2024), SVHN (Pradhan et al. 2024),and CIFAR-10 (Aslam and Nassif 2023) are utilized to make performance evaluation.
  3. 3.
    Training Setup:
    • Training Parameters. We utilize a CNN training architecture. The batch size is 32, the local update round $e$ is set to 5, and the global training round $E$ is set to 100. The learning rate $\eta$ is configured to 0.01 and the SGD momentum is set to 0.05.
  4. 4.
@@
    • TRUST. In (Wang et al. 2023), the authors introduce a trust-age of information (AoI)-aware co-design scheme (TACS), employing a learning-based trust model and trust-AoI-aware scheduling to optimize data selection for plant control dynamically.
@@
    • Test accuracy. Test accuracy is a crucial metric for model training in FL, representing the performance that model owners can achieve with their trained models.
@@
Figure 2: The prediction accuracy and training loss with 10% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
Figure 3: The prediction accuracy with 10%,30%,50% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
@@
In Figure 2, we analyze the variations in test accuracy and training loss over multiple training rounds for four distinct mechanisms, specifically under conditions where only 10% of clients are classified as low quality. Our proposed mechanism stands out by achieving the highest performance across four real datasets. This success can be attributed to the effective integration of our AHSMM and a greedy algorithm, which work synergistically to accurately predict fluctuations in client learning and communication quality. By optimizing the participation of low-quality clients, our approach significantly enhances overall training outcomes. In contrast, the TRUST mechanism, despite employing a predictive approach, lacks an effective client distribution strategy. This deficiency leads to suboptimal performance, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism incorporates the AHSMM to forecast client behavior, but it does not allocate clients efficiently, leading to less effective training sessions. Although the GUROBI mechanism can determine the optimal client allocation scheme, it does not incorporate client quality predictions, which hampers its ability to promptly exclude low-quality clients from participating in training, ultimately affecting training efficiency. Overall, our mechanism demonstrates a superior ability to navigate the complexities of client quality by leveraging predictive modeling and strategic allocation, ensuring robust training performance in Federated Learning environments.
@@
In Figure 3, we analyze the comparative performance of four mechanisms across scenarios characterized by varying proportions of low-quality clients. As the percentage of low-quality clients increases, all mechanisms demonstrate a decline in both training and testing accuracy, albeit to different extents. Notably, our proposed mechanism consistently delivers the best results across four real datasets, showcasing its robustness in challenging environments. The superior performance of our mechanism can be attributed to the integration of the AHSMM and a greedy-based client allocation scheme. This combination effectively predicts fluctuations in client learning and communication quality, enabling efficient client allocation that significantly enhances model training quality, even under adverse conditions. By optimizing the participation of higher-quality clients, we mitigate the negative impact of low-quality clients on overall performance. In contrast, while the TRUST mechanism is capable of identifying unreliable clients, it lacks an efficient client distribution strategy. This limitation results in poorer training outcomes compared to our mechanism, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism employs the AHSMM to accurately predict changes in client training quality but relies on a random allocation strategy during the client distribution phase. Consequently, this randomness undermines the final training effectiveness, leading to suboptimal results. The GUROBI mechanism, despite its potential for determining optimal client distributions, performs the worst in our experiments. Its inability to accurately predict changes in client learning quality restricts its effectiveness, as it cannot exclude low-quality clients in a timely manner. Our findings reveal a notable 8.7% increase in test accuracy and a 15.3% reduction in training loss compared to existing baselines, highlighting the efficacy of our approach in enhancing model performance in Federated Learning settings. These results underscore the importance of both predictive modeling and strategic client allocation in achieving high-quality training outcomes.
@@ Conclusion
In this work, we proposed TRAIL, a novel mechanism designed to address the dynamic challenges in SD-FL. TRAIL significantly improves training efficiency and robustness by integrating an AHSMM to predict client states and contributions and a greedy algorithm to optimize client-server associations. Experimental results on real-world datasets show an improvement compared to state-of-the-art baselines.
For future work, we aim to enhance TRAIL by incorporating more complex dynamic modeling, multi-objective optimization, robustness against failures, scalability for large-scale systems, and stronger privacy and security mechanisms. These improvements will further broaden TRAIL’s applicability and effectiveness in real-world decentralized learning scenarios.

### v3

@@ 8             $x_{i}=1$;
From Theorem 1, the global training loss minimization initially outlined in the problem can be reinterpreted to focus primarily on minimizing the parameter $B$. This revised formulation, therefore, positions $B$ as the central target for reduction, aiming to directly influence and improve the overall system performance by addressing the underlying factors contributing to $B$ ’s value, i.e.,
@@
To address this nonlinear integer programming problem, we propose a greedy algorithm outlined in Algorithm 1, with a time complexity of $O(nm\log nm)$.
@@
This section evaluates our proposed mechanism using real-world datasets to demonstrate its effectiveness and practicality. We begin by introducing the experiment setting. Then, we present our experimental comparisons’ results, highlighting our mechanism’s performance relative to the baselines.
@@
We provide a detailed explanation of the fundamental experimental setup, including the basic setup, datasets, training configurations, baselines, and evaluation metrics.
Basic Setup: We design a SD-FL system comprising five edge servers and fifty clients, with each client assigned 1,000 local training samples. To emulate real-world challenges, 10%, 30%, and 50% of the clients gradually experience degradation in both training quality (e.g., training accuracy) and communication quality (e.g., packet loss) as training progresses.
Datasets: Real-world Datasets. Four standard real-world datasets, e.g. MNIST (Rana, Kabir, and Sobur 2023), EMNIST (Majeed et al. 2024), SVHN (Pradhan et al. 2024),and CIFAR-10 (Aslam and Nassif 2023) are utilized to make performance evaluation.
Training Configurations: Training Parameters. We adopt a CNN architecture for its effectiveness in image processing tasks. The batch size is 32, balancing computational efficiency and model performance. Each client performs 100 local training rounds ($T_{1}=100$) before aggregation, with 100 inter-cluster aggregations ($T_{2}=100$) to synchronize updates across edge servers. The learning rate ($\eta$) is set to 0.01, ensuring stable and efficient optimization, and SGD with a momentum of 0.05. The model uses ReLU as the activation function for non-linearity and cross-entropy loss for classification tasks.
@@
  • TRUST. In (Wang et al. 2024c), the authors introduce a trust-age of information (AoI)-aware co-design scheme (TACS), employing a learning-based trust model and trust-AoI-aware scheduling to optimize data selection for plant control dynamically.
@@
  • Test accuracy. Test accuracy measures a model’s performance on unseen data, reflecting its generalization ability and effectiveness in SD-FL, critical for real-world usability and reliability.
@@
Figure 2: The test accuracy and training loss in scenarios with 10% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
Figure 3: The test accuracy in scenarios with 10%,30%,50% low-quality clients: (a) MNIST, (b) EMNIST, (c) CIFAR10, and (d) SVHN.
@@
In Figure 2, we analyze the variations in test accuracy and training loss over multiple training rounds for four distinct mechanisms, specifically under conditions where only 10% of clients are classified as low quality. Our proposed mechanism stands out by achieving the highest performance across four real datasets. This result is due to the effective integration of our AHSMM model with a greedy algorithm, which works together to predict fluctuations in client learning and communication quality reasonably. By optimizing the participation of low-quality clients, our mechanism significantly enhances overall training outcomes. In contrast, the TRUST mechanism, despite employing a predictive approach, lacks an effective client distribution strategy. This deficiency leads to suboptimal performance, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism incorporates the AHSMM to forecast client behavior, but it does not allocate clients efficiently, leading to less effective training sessions. Although the GUROBI mechanism can determine the optimal client allocation scheme, it does not incorporate client quality predictions, which hampers its ability to promptly exclude low-quality clients from participating in training, ultimately affecting training efficiency. Overall, our mechanism demonstrates a superior ability to navigate the complexities of client quality by leveraging predictive modeling and strategic allocation, ensuring robust training performance in SD-FL environments.
@@
In Figure 3, we analyze the comparative performance of four mechanisms across scenarios characterized by varying proportions of low-quality clients. As the percentage of low-quality clients increases, all mechanisms demonstrate a decline in both training and testing accuracy, albeit to different extents. Notably, our proposed mechanism consistently delivers the best results across four real datasets, showcasing its robustness in challenging environments. The superior performance of our mechanism can be attributed to the integration of the AHSMM and a greedy-based client allocation algorithm. This combination effectively predicts fluctuations in client learning and communication quality, enabling efficient client allocation that significantly enhances model training quality, even under adverse conditions. By optimizing the participation of higher-quality clients, we mitigate the negative impact of low-quality clients on overall performance. In contrast, while the TRUST mechanism is capable of identifying unreliable clients, it lacks an efficient client distribution strategy. This limitation results in poorer training outcomes compared to our mechanism, as it fails to adaptively manage client participation based on their quality. Similarly, the RANDOM mechanism employs the AHSMM to accurately predict changes in client training quality but relies on a random allocation strategy during the client distribution phase. Consequently, this randomness undermines the final training effectiveness, leading to suboptimal results. The GUROBI mechanism, despite its potential for determining optimal client distributions, performs the worst in our experiments. Its inability to accurately predict changes in client learning quality restricts its effectiveness, as it cannot exclude low-quality clients in a timely manner. Experiments reveal a notable 8.7% increase in test accuracy and a 15.3% reduction in training loss compared to existing baselines, demonstrating the superiority of our mechanism in SD-FL settings. These results underscore the importance of both predictive modeling and strategic client allocation in achieving high-quality training outcomes.
@@ Conclusion
This paper proposes TRAIL, a novel mechanism designed to address the dynamic challenges in SD-FL. TRAIL integrates an AHSMM to accurately predict client states and contributions and a greedy algorithm to optimize client-server associations, effectively minimizing global training loss. Through convergence analysis, the impact of client-server relationships on model convergence is theoretically assessed. Extensive experiments conducted on four real-world datasets demonstrate that TRAIL improves test accuracy and training loss, significantly outperforming state-of-the-art baselines. This work highlights the potential of combining predictive modeling and strategic client allocation to enhance efficiency, robustness, and performance in distributed learning systems.
@@ Acknowledgments
Appendix
Proof of Theorem 1
Assumption 1: Strong Convexity and Lipschitz Continuity Assume the global loss function $\mathrm{F}(g)$ is $\mu$-strongly convex and its gradient is $L$-Lipschitz continuous. This implies that for any $g$ and $g^{\prime}$:
  • Strong convexity:
    $F\left(g^{\prime}\right)\geq F(g)+\left\langle\nabla F(g),g^{\prime}-g\right\rangle+\frac{\mu}{2}\left\|g^{\prime}-g\right\|^{2}.$
  • Gradient Lipschitz continuity:
    $\left\|\nabla F(g)-\nabla F\left(g^{\prime}\right)\right\|\leq L\left\|g-g^{\prime}\right\|.$
Assumption 2: Assume that during training, randomness is introduced (e.g., due to random client participation), and these random factors are independent and identically distributed (i.i.d.).
Definitions: Let $g_{t}$ denote the global model at iteration $t$ , and $g^{*}$ denote the global optimal model. Let $omega_{1}$ and $\omega_{2}$ be constants related to the system, and $B$ is a cumulative error term. In FL, the global model update can be expressed as:
$g_{t+1}=g_{t}-\lambda\nabla F\left(g_{t}\right)+\eta_{t},$
where $\eta_{t}$ represents the noise or error term due to factors like client sampling and unreliable communication. Since $\lambda=\frac{1}{L}$, we have:
$g_{t+1}=g_{t}-\frac{1}{L}\nabla F\left(g_{t}\right)+\eta_{t}.$
Consider the difference:
$F\left(g_{t+1}\right)-F\left(g^{*}\right).$
Using the $L$-Lipschitz continuity of the gradient, we have:
$F\left(g_{t+1}\right)\leq F\left(g_{t}\right)+\left\langle\nabla F\left(g_{t}\right),g_{t+1}-g_{t}\right\rangle+\frac{L}{2}\left\|g_{t+1}-g_{t}\right\|^{2}$
Substituting $9_{\{}t+1\}-g_{t}$, we get:
$g_{t+1}-g_{t}=-\frac{1}{L}\nabla F\left(g_{t}\right)+\eta_{t},$
so
$\left\|g_{t+1}-g_{t}\right\|\leq\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|.$
Substitute back into the inequality:
$F\left(g_{t+1}\right)\leq F\left(g_{t}\right)-\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|^{2}+\left\langle\nabla F\left(g_{t}\right),\eta_{t}\right\rangle$
$+\frac{L}{2}\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}.$
Take the expectation of both sides. Assuming that $\eta_{t}$ has zero mean (the noise is mean zero), we have:
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{1}{L}E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$\frac{L}{2}E\left[\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}\right].$
Expanding the squared term:
$\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|+\left\|\eta_{t}\right\|\right)^{2}=\left(\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|\right)^{2}+$
$2\frac{1}{L}\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|+\left\|\eta_{t}\right\|^{2}.$
Therefore, we can get
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-$
$\left(\frac{1}{L}-\frac{1}{2L}\right)E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right].$
Note that $\frac{1}{L}-\frac{1}{2L}=\frac{1}{2L}$. Thus,
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{1}{2L}E\left[\left\|\nabla F\left(g_{t}\right)\right\|^{2}\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right].$
Since $F(g)$ is $\mu$-strongly convex, we have:
$\left\|\nabla F\left(g_{t}\right)\right\|^{2}\geq 2\mu\left(F\left(g_{t}\right)-F\left(g^{*}\right)\right)$
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-\frac{\mu}{L}E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+$
$LE\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]+\frac{L}{2}E\left[\left\|\eta_{t}\right\|^{2}\right.$
Introduce constants $\omega_{1}$ and $\omega_{2}$ satisfying:
$E\left[\left\|\nabla F\left(g_{t}\right)\right\|\left\|\eta_{t}\right\|\right]\leq\omega_{1}E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\omega_{2}B,$
and
$E\left[\left\|\eta_{t}\right\|^{2}\right]\leq B$
where $B$ is the cumulative error term defined earlier. Substitute these into the inequality:
$E\left[F\left(g_{t+1}\right)\right]\leq E\left[F\left(g_{t}\right)\right]-$
$\left(\frac{\mu}{L}-L\omega_{1}\right)E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\left(L\omega_{2}+\frac{L}{2}\right)B.$
Thus,
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq$
$\left(1-\frac{\mu}{L}+L\omega_{1}\right)E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\left(L\omega_{2}+\frac{L}{2}\right)B.$
Note that $D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}$. We can adjust constants $\omega_{1}$ and $\omega_{2}$ (since they are related to $B$) to satisfy:
$L\omega_{1}=\frac{4\omega_{2}\mu B}{L}.$
Therefore, the inequality becomes:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq DE\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}.$
By iteratively applying the inequality from $t=0$ to $t$, we get:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq D^{t+1}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+$
$\frac{2\omega_{1}B}{L}\sum_{k=0}^{t}D^{k}.$
Using the geometric series formula:
$\sum_{k=0}^{t}D^{k}=\frac{1-D^{t+1}}{1-D}$
we have:
$E\left[F\left(g_{t+1}\right)-F\left(g^{*}\right)\right]\leq D^{t+1}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+$
$\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t+1}}{1-D}.$
To align with the statement of Theorem 1, we replace $t+1$ with $t$, so the final result is:
$E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]\leq D^{t}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t}}{1-D}.$
We have thus proven that for the learning rate $\lambda=\frac{1}{L}$, the expected difference in function values satisfies:
$E\left[F\left(g_{t}\right)-F\left(g^{*}\right)\right]\leq D^{t}E\left[F\left(g_{0}\right)-F\left(g^{*}\right)\right]+\frac{2\omega_{1}B}{L}\cdot\frac{1-D^{t}}{1-D},$
where
$D=1-\frac{\mu}{L}+\frac{4\omega_{2}\mu B}{L}.$
In the above proof, $B$ is the cumulative error term, which is related to factors like client selection and communication quality. Specifically, $B$ is defined as:
$B=\sum_{m\in\mathcal{S}}\frac{1}{N_{m}^{(S)}}\Psi$
where
$\Psi=\left[\sum_{i\in\mathcal{U}}n_{i}d_{i,m}\left(D_{m}-1+I\left(TL_{i,s}<\Theta\right)\right)\right],$
$N^{(S)}_{m}$ is the total data size of clients assigned to server $m$:
$N_{m}^{(S)}=\sum_{i\in\mathcal{U}}n_{i}d_{i,m},$
$D_{m}$ is a distribution parameter:
$D_{m}=\frac{1}{1+|\mathcal{S}|},$
and $\mathbb{I}(TL_{i}<\Theta)$ is an indicator function that equals 1 if the trust level of client $i$ is below the threshold $\Theta$, and 0 otherwise.
By minimizing $B$, we can directly influence and improve the overall performance of the system. Through the detailed derivation above, we have proven Theorem 1. In the proof, we have made standard assumptions such as the strong convexity of the loss function and the Lipschitz continuity of the gradient. We have also taken into account the randomness and error terms during the training process, ultimately deriving an upper bound on the expected difference in the global loss function’s value.